# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

### v1

@@ Abstract
The Mountain pine beetle (MPB) is a destructive forest pest that undergoes boom-bust cycles every 20-40 years. It also has unusual life-history traits, including a chemosensory affinity for large pine trees, coordinated attacks facilitated by aggregation and anti-aggregation pheromones, and variable dispersal that increases in the absence of trees. Using an empirically calibrated model, we show that MPB’s distinctive life-history characteristics are responsible for its distinctive population dynamics. Specifically, three life-history traits are necessary for episodic boom-bust dynamics: density-dependent dispersal, where beetles disperse once most of the large trees have been killed; an Allee effect, which requires a threshold number of beetles to overcome tree defenses; and a short generation time, notably in comparison with their pine tree hosts. In addition to explaining the qualitative behavior of MPB dynamics, these three ingredients explain residual tree density, the duration of outbreaks, and the average waiting time between outbreaks. The peak number of beetles, believed to be the primary factor driving range expansion, is mostly a consequence of MPB’s constitutive high fecundity. However, two life history traits — MPB’s size-dependent fecundity and preference for large trees — are responsible for 25% of the peak number of beetles. By identifying patterns across the extensive MPB literature, and by integrating data across multiple sources, we develop a highly accurate description of MPB’s density-dependent interactions with lodgepole pine. We conclude with a simplified mechanistic model, distilled to two difference equations, effectively capturing the essence of MPB outbreak dynamics.
Keywords: mountain pine beetle, outbreak, excitable, density dependence
@@ 1 Introduction
Over the last 25 years, the mountain pine beetle (MPB; Dendroctonus ponderosae) has killed over 140,000 $\text{km}^{2}$ of pine forests (Fig. 1). Populations of MPB grew synchronously, from Arizona to British Columbia, coalescing to form the largest insect outbreak ever recorded (Taylor et al.,, 2006). By comparison, the previous continent-scale outbreak of MPB in the 1980s appears minuscule. On the local scale (around $1\text{km}^{2}$), MPB populations grow geometrically, plateau in abundance for 1 to 2 years, and then rapidly decline (Fig. D.3).
Figure 1: Time series of MPB damage, as measured by contiguous area of tree crown mortality. To estimate contiguous tree crown mortality, we took fixed-wing Aerial Overview Survey data from both the United States and Canada, used histogram matching to harmonize across different data modalities, and then using the “FHP method” of Egan et al., (2020). The All regions category includes British Columbia (B.C.), Alberta, and Regions 1–6 as defined by the United States Forest Service; 1997 is the first year that all regions were surveyed. The similarity of the All regions and B.C. time series reveals that most of the damage was concentrated in British Columbia.
@@
Figure 2: Progression of the 2000s MPB outbreak, also known as the hyperepidemic. The data comes from fixed-wing Aerial Overview Surveys (United States and British Columbia) and Heli-GPS surveys (Alberta).
@@
Why are mountain pine beetles like this? Why are they so destructive, why are outbreaks widespread, and why are outbreaks episodic with rapid booms and busts? There are several ways to answer these questions, and the relevance of each answer will depend on the underlying point of comparison. For instance, if we ask why the 2000s outbreak was large, relative to past outbreaks, then climate change provides an answer (Carroll et al., 2006a, ; Alfaro et al.,, 2009; Creeden et al.,, 2014). Widespread drought conditions caused synchronized increases in tree susceptibility. At the same time, warmer winter temperatures reduced overwintering mortality in beetle larvae. Fire suppression has also played a role, by enabling the buildup of large-diameter pines (Taylor et al.,, 2006). Alternatively, we might wonder why MPB is so destructive, relative to other insect pests. Most bark beetles are secondary bark beetles, meaning that they infest downed, diseased, or otherwise weakened trees. By contrast, MPB is a primary bark beetle, which means that it attacks and kills healthy host trees, thus accessing all the resources of healthy phloem. It is notable that of the four main bark beetles in British Columbia, the two most destructive (mountain pine beetle and the western balsam bark beetle) are obligate primary beetles, whereas the two least destructive (spruce beetle and Douglas-fir beetle) usually prefer downed trees. A comparative analysis of bark beetles found that the occurrence of outbreaks mostly depends on the availability of susceptible host trees, not life-history traits (Koricheva et al.,, 2012, though the primary/secondary bark beetle distinction was not included in the analysis). Anecdotally, this conclusion is supported by the fact that MPB’s primary host, lodgepole pine, is the most abundant tree in British Columbia, Finally, we might ask why MPB was so much more destructive in British Columbia, relative to the western United States (see Figs.2 & 2). Although this spatial pattern is not well-researched, we suspect that it involves the high density and spatial continuity of lodgepole pine in the Chilcotin Plateau –– the epicenter of the last two outbreaks in British Columbia.
In this paper, we explore how the distinctive life history of the mountain pine beetle affects its population dynamics. We ask: Why do mountain pine beetles experience episodic boom-bust dynamics, relative to an archetypical predator-prey system in which limit cycles or equilibria often appear (e.g., Rosenzweig and MacArthur,, 1963)? To answer this question we 1) build and validate a realistic model of MPB dynamics, and 2) use the model to perform a causal analysis (Fig. 3). The causal analysis involves counterfactual simulations of the model where the traits are turned on/off, followed by a comparison of simulated dynamics and real outbreaks.
Figure 3: Graphical representation of our inferential approach. First, we develop and validate a realistic model of MPB dynamics. Second, we perform a causal analysis in which we turn life-history traits on or off. In the example of the causal analysis, the red crosses respectively indicate that life-history trait # 1 has been off, and that the simulated system does not match reality with respect to Metric #2 (e.g., outbreak periodicity). The dashed red arrow indicates that life-history trait #1 is responsible for realistic values of Metric #2.
@@ 1.1 The biology of mountain pine beetle and lodgepole pine
For additional context, we provide a brief natural history of the MPB/lodgepole pine system. More detailed examinations are provided by Safranyik and Carroll, (2006), Raffa, (2001), and Lotan et al., (1983). The univoltine life cycle of MPB begins with adult beetles emerging from their natal trees in July or August. After emergence these beetles fly in search of new host trees, typically targeting large pines. Suitable host species include lodgepole, ponderosa, limber, whitebark, and western white pines. MPB has infested Jack pine and Rocky Mountain bristlecone pine, but the general suitability of these species is a matter of ongoing research (Bentz,, 2019; Bentz et al.,, 2022; Bleiker et al.,, 2023). Individual beetles may fly around for several days, but few beetles survive more than 5–10 days outside of the tree (Reid and Reid,, 2008; Fuchs and Borden,, 1985). Emergence from natal trees is synchronous (due to temperature-dependent development thresholds; Powell and Bentz,, 2009; Logan and Powell,, 2001), but the synchronicity is not perfect, and so the flight period of the entire population may last up to 30 days.
@@
Because individual beetles cannot be tracked, the dispersal phase of the life cycle remains somewhat enigmatic. However, available evidence points to three modes of dispersal:short-, medium-, and long-range dispersal. First, a large majority of beetles disperse less than 30-50 meters under the canopy (Safranyik et al.,, 1992; Robertson et al.,, 2007; for the Douglas-fir beetle Dodds and Ross,, 2002). Second, a minority of beetles disperse up to 5 km (Robertson et al.,, 2009; Powell and Bentz,, 2014; Howe et al.,, 2021; Koch et al.,, 2021), performing pioneer attacks and aggregating to form semi-regular patterns of “spot infestations” (clusters of infested trees) across the landscape (Strohm et al.,, 2013). Finally, 0–2.5% of beetles fly above the canopy (Robertson et al.,, 2009; Safranyik et al.,, 1992), though experts have speculated that the percentage may be higher in stands where most suitable host trees have already been killed (Bleiker,, 2019). These above-canopy beetles may get swept away and carried up to 300 km by upper atmospheric winds (Hiratsuka et al.,, 1982; Cerezke,, 1989; Jackson et al.,, 2008).
@@
Upon locating a suitable host, MPB females initiate attacks. The trees fight back by exuding resin, which serves to entrap or pitch out (i.e. expel) attacking MPB. Female beetles release aggregation pheromones that attract additional beetles. This strategy, employed by many bark beetle species, is called a mass attack. Depending on the beetle population density, tree desirability, and the tree’s defensive capability, successful mass attacks may last from a couple hours to a week, continuing until the tree’s resin supply is depleted. Subsequently, females bore into the tree and begin excavating egg galleries. Mating occurs within the tree, and males will either seek out additional females or “settle down” to help with gallery defense.
Eggs hatch within one to two weeks, and the newly hatched larvae immediately start feeding on the tree’s phloem. Tree death is hastened by MPB’s fungal symbionts, blue stain fungi spp., which colonize the tree’s transport vessels. The beetles carry fungal spores in specialized mouthparts, thereby spreading and inoculating the fungi. In return, the fungi kill the host trees, thus preventing MPB larvae from being killed by future resinosis, were the tree to recover from the initial mass attack (Safranyik et al.,, 1975). Spores that line pupal chambers also serve as a source of nutrition for pre-emergence adult beetles (Reid,, 1958)
In late fall, the larvae enter a dormant state. Despite the accumulation of antifreeze chemicals in their tissue, late-instar larvae face appreciable mortality when temperatures are less than -30°C, and near-100% mortality when temperatures are less than -40°C (Wygant,, 1940; Safranyik and Linton,, 1991; Logan et al.,, 1995). Less extreme temperatures can kill the less resistant life-stages (i.e., early-instar larvae and pupae), particularly during cold snap in the fall or spring (Régnière and Bentz,, 2007; Rosenberger et al.,, 2017). Come spring, surviving larvae resume feeding and pupate in the late spring or early summer. Adults emerge in the summer, completing the univoltine life cycle. It is around this time, approximately one year after the initial attack, that the foliage of all infested trees has turned a telltale rust-red color. Within 2–4 years after the initial attack, most needles will have dropped from the tree (Klutsch et al.,, 2009; Audley et al.,, 2021).
@@
Lodgepole pine and ponderosa pine are the predominant hosts of MPB, simply because they are the most abundant pine species in MPB’s geographic range and elevational niche. Lodgepole pine dominates in British Columbia, western Alberta, and the American Rocky Mountain states, whereas ponderosa pine is more abundant in the American southwest and the Sierra mountains (Beaudoin et al.,, 2014; Ellenwood et al.,, 2015). Both species are shade-intolerant and fire-resistant. They tend to be seral (i.e., persisting as a non-climax species in a fire cycle), taking advantage of high-light conditions following a stand-leveling fire (Fryer,, 2018; Anderson,, 2003; Lotan et al.,, 1985). Because MPB rarely kills all of the mature trees in a stand, and because pines produce many cones, no pine species faces recruitment limitation during recovery from MPB outbreaks. This is doubly true for lodgepole pine, which has closed cones with seeds that remain viable for decades, effectively creating a seed bank (Teste et al., 2011a, ; Teste et al., 2011b, ). Due to its abundant seeds and its ability to thrive in high-light conditions, lodgepole pine can take advantage of the light gaps created in the years following a mountain pine beetle attack, resulting in the recruitment of seedlings into approximately even-aged cohorts (Axelson et al.,, 2010)
@@
During epidemics, MPB seeks out large healthy trees with abundant phloem resources. However, in their low-density endemic state, there are not enough beetles to mass attack a single healthy tree, and thus there is a collective behavioral switch where beetles seek out trees that are weakened or stressed, typically due to root disease, wind damage, or lightning (Tkacz and Schmitz,, 1986; Carroll et al., 2006b, ). Endemic populations tend to remain at low density, primarily due to resource limitation, but also because of parasitoids, woodpeckers, and competition with secondary bark beetle species (Amman,, 1983; Safranyik and Carroll,, 2006). Endemic populations of mountain pine beetle (MPB) can transition into an incipient-epidemic state when they successfully breed within trees that have thick phloem. This scenario can occur if a healthy tree experiences acute stress (such as wind damage or lightning) right before the MPB flight period, such that secondary bark beetles have not yet extensively colonized the tree (Bleiker et al.,, 2014). Such stressors can trigger local outbreaks (on the scale of square kilometers). However, widespread outbreaks are usually preceded by widespread drought conditions, which synchronously weaken the resin defenses of many trees (Alfaro et al.,, 2009; Creeden et al.,, 2014).
1.2 Mountain pine beetle life history traits

### v2

# Explaining excitable population dynamics in bark beetles: From life history to large, episodic outbreaks

@@ Abstract
Bark beetles are significant forest pests, with some species capable of causing widespread tree mortality. Among these, the mountain pine beetle (MPB) stands out for its exceptionally destructive outbreak in the 2000s. We use MPB as a case study to explore the concept of excitable dynamics, where ephemeral perturbations produce large excursions from equilibrium. Our empirically-calibrated model of reveals five features of MPB biology: stand density-dependent dispersal, an Allee effect, time-scale separation between beetle and tree life cycles, tree size-dependent fecundity, and a large-tree preference. The first three features explain MPB’s characteristic boom-bust dynamics, while the latter two explain outbreak magnitude. Other bark beetles lack one or more of these traits, partially explaining their generally lower impact. However, predicting bark beetle impact requires consideration of both life history and landscape factors: total damage increases linearly with host-tree biomass, but this relationship holds only for irruptive (i.e., excitable) beetle species. We distill our findings into a minimal mechanistic model that captures the essence of irruptive bark beetle dynamics. This model firmly establishes MPB as one of the first empirical examples of excitable dynamics in ecology.
Keywords: bark beetle, mountain pine beetle, outbreak, excitable dynamics, lodgepole pine, forest pest, population dynamics, density dependence
@@ 1 Introduction
Bark beetle (Coleoptera: Curculionidae, Scolytinae) outbreaks occur worldwide (Lantschner and Corley,, 2023), with recent epidemics in western North America causing widespread tree mortality (Meddens et al.,, 2012; Taylor et al.,, 2006). These outbreaks exhibit a distinctive pattern: unpredictable and explosive population growth followed by rapid declines (Fig. E.1). The episodic “boom-bust” pattern contrasts sharply with the population dynamics of most species, which typically show either stochastic fluctuations around a weakly-attracting equilibrium or cycles driven by resource-consumer interactions (Louca and Doebeli,, 2015; Hassell et al.,, 1976; Ellner and Turchin,, 1995). In this paper, we demonstrate that certain bark beetle species are unique in that their population dynamics can be classified as “excitable”.
Excitable population dynamics are characterized by several features that set them apart from other types of dynamic behaviors (Murray,, 1989; Morozov and Petrovskii,, 2009; Hastings et al.,, 2021). Perturbations trigger large excursions from equilibrium, leading to dramatic population increases from minor or transient environmental changes. These excursions involve both positive and negative feedback mechanisms, creating a self-reinforcing cycle of growth followed by an eventual crash. Excitable systems often exhibit “slow-fast dynamics” (Hastings et al.,, 2018), where one state variable changes rapidly (e.g., beetle population size), while another changes slowly (e.g., tree population size). This results in a refractory period, the minimum time between excursions. In summary, the notion of excitability (while lacking a formal mathematical definition; Strogatz,, 1994, pg. 118) involves perturbations causing large but ultimately transient deviations from a stable equilibrium. The canonical example of excitable dynamics is the action potential of a neuron (Hodgkin and Huxley,, 1952): A perturabation (neurotransmitters from neighboring neurons) and positive feedback causes a neuron to fire, and slow-fast dynamics (related to conformal changes in ion channels) result in a refractory period. In the case of neurons, excitability is adaptive, ensuring unidirectional signal propagation and regulating firing frequency (Johnston and Wu,, 1994).
Bark beetle outbreaks display these hallmarks of excitability. A perturbation such as drought or wind damage can increase tree susceptibility to beetle attack (Singh et al.,, 2024). This initiates a positive feedback loop: larger beetle populations overcome bigger, better-defended trees, further increasing their numbers. Irruptive bark beetles must mass attack host trees to overcome toxic resin defenses (Thalenhorst,, 1958; Raffa and Berryman,, 1983; Boone et al.,, 2011), using aggregation pheromones and tree volatiles to focus their efforts (Raffa,, 2001; Seybold et al.,, 2006). Many bark beetle populations are thought to exhibit a strong Allee effect, wherein per capita growth rates become negative at sufficiently low population densities (Taylor,, 2012). Negative feedback when beetles run out of suitable or detectable host trees. Post-outbreak, trees must grow until their phloem or cambium is thick enough to support beetle development, which generates a slow-fast dynamic and a minimum interval between outbreaks. The perturbation that triggers the outbreak, however, is effectively random — there is no maximum interval. Therefore, bark beetle outbreaks occur irregularly (episodic) rather than on a fixed schedule (cyclical).
We will demonstrate that bark beetles are unique in ecology as the only known taxonomic group known to exhibit excitable population dynamics. While the spruce budworm (Choristoneura fumiferana) is sometimes cited as an example, this is probably incorrect. An influential mathematical paper (Ludwig et al.,, 1978) proposed that excitability in the spruce budworm system could result from interactions between trees, spruce budworms, and their avian predators. Specifically, birds with a type III functional response (Holling,, 1959) led to low predation rates at low budworm density due to prey-switching (preventing budworm extirpation between outbreaks), and saturating rates at high density due to handling time constraints (generating positive density dependence during incipient outbreaks). The slow part of the slow-fast dynamics is attributed to the fact forest quality rebounds slowly, and that avian predation is assumed to be proportional to forest quality. However, current evidence indicates that spruce budworms undergo noisy cycles due to their interactions with parasitoids (Royama,, 1984; Berryman,, 1996; Royama,, 1997; Régnière and Nealis,, 2007; Pureswaran et al.,, 2016) rather than excitable dynamics. Although budworms may face mate-finding challenges at low densities (Régnière et al.,, 2013), there is little evidence for a demographic Allee effect over larger spatial scales.
Examples of excitable dynamics in ecology have been explored theoretically but lack empirical support. Models of algal blooms suggest the possibility of excitable dynamics triggered by a temporary influx of nutrients, with positive feedback arising from either top-down or bottom-up control. The top-down mechanism is a saturating functional response where zooplankton become satiated (Truscott and Brindley, 1994a, ; Truscott and Brindley, 1994b, ); negative feedback comes from predation, and the slow-fast dynamics result from disparate generation times of zooplankton and phytoplankton. The bottom-up mechanisms involves the formation of algal mats, which improves nutrient cycling and accelerate algal growth (Huppert et al.,, 2002, 2004); the negative feedback comes from nutrient depletion, and the slow-fast dynamics result from slow replenishment of nutrients from runoff, decomposition, etc. Excitable algal blooms are likely to occur somewhere, given the diverse character of algal blooms worldwide (Glibert et al.,, 2018), anecdotes of threshold-effects (Pearsall,, 1932; Lund,, 1950), and correlations between the bloom occurrence and environmental fluctuations (e.g., Freund et al.,, 2006). To the best of our knowledge, however, there is no well-justified and empirically calibrated model that provides clear evidence for excitable algal blooms. Small mammal populations can experience severe outbreaks, but these are typically classified as pulse-driven rather than irruptive (Andreassen et al.,, 2021). Pulse-driven outbreaks occur due to significant increases in resource availability, such as abundant vegetation following unusually high precipitation, rather than small or transient perturbations characteristic of excitable systems (Holt,, 2008).
Some bark beetle species undergo pulse-driven outbreaks triggered by logging, windfall, or drought. Definitionally, these outbreaks subside shortly after the disturbance ends. For example, prolonged drought in the western United States from 1995–2004 (Breshears et al.,, 2005) depleted trees’ defensive resin reserves, which led to an outbreak of pinyon Ips beetles. The outbreak peaked from 2002 to 2004, and infestation densities fell by nearly two orders of magnitude after the drought ended (Raffa et al.,, 2008). In contrast, excitable bark beetle outbreaks persist long after the initial disturbance subsides. An outbreak of the North American Spruce Beetle (Dendroctonus rufipennis Kirby) lasted 13 years following a single windthrow event in 1939 (Massey and Wygant,, 1954; Anderson et al.,, 2010; Forest Health Protection,, 2011) and an outbreak of the mountain pine beetle (MPB; Dendroctonus ponderosae Hopkins) continuing for 10–15 years after the initiating drought ended in 2000 (Creeden et al.,, 2014). Bark beetles that exhibit excitable dynamics are termed irruptive or aggressive, while those with pulse-driven outbreaks are classified as semi-aggressive or opportunistically aggressive (Lantschner and Corley,, 2023; Weed et al.,, 2015).
In this paper, we ask: why do some bark beetles species display excitable population dynamics, and why are some outbreaks so large? To answer these questions, we examine the mountain pine beetle as a case study, using insights from this species and comparative analysis to better understand population patterns across bark beetle species. Our methodology involves four key steps:
  1. 1.
    Developing and validating a realistic model of mountain pine beetle dynamics, incorporating known biological features.
  2. 2.
    Performing a causal analysis through counterfactual simulations, manipulating various biological features to determine their impact on population dynamics.
  3. 3.
    Generalizing the results to other irruptive bark beetle species based on shared life history characteristics.
  4. 4.
    Building a minimal mechanistic model which contains only the necessary processes for excitability
Additionally, we explore the relationship between excitability and the magnitude of bark beetle outbreaks (i.e., the peak number of beetles). In theory, the two concepts are orthogonal — excitability only implies that outbreaks are large relative to the initiating perturbation. In practice, however, the two concepts are connected. Later, we show that the total area affected by a beetle species is proportional to the combined biomass of their host trees, but that this relationship only applies to excitable (also known as irruptive) species; the relationship does not hold for beetles with pulse-driven outbreaks.
The mountain pine beetle serves as our model species, chosen for its well-documented life history and its significant impact. The most recent outbreak of MPB (ca. 2000–2015) stands as the largest recorded insect outbreak in history (Taylor et al.,, 2006), killing up to 8.5 million hectares of tree crowns across western North America (Meddens et al.,, 2012), and killing $>50\%$ of merchantable pine volume in British Columbia (British Columbia Government,, 2016). This event had far-reaching consequences, including transforming British Columbia’s forests from a carbon sink to a source (Kurz et al.,, 2008), inflicting billions of dollars in economic damage (Corbett et al.,, 2016), and altering hydrological functioning (Schnorbus,, 2011).
@@
Figure 1: Progression of the latest large-scale mountain pine beetle outbreak. The data comes from fixed-wing Aerial Overview Surveys (United States and British Columbia) and Heli-GPS surveys (Alberta). Pine forest data comes from Beaudoin et al., (2014) and (Ellenwood et al.,, 2015).
@@
Of course, the question of “Why is mountain pine beetle so destructive?” has been studied extensively. Climate change is the principal reason for the severity of the outbreak in the 2000s relative to past outbreaks (Carroll et al., 2006a, ; Alfaro et al.,, 2009; Creeden et al.,, 2014). Widespread drought conditions caused synchronized increases in tree susceptibility. At the same time, warmer winter temperatures reduced overwintering mortality in beetle larvae. Fire suppression has also played a role, by enabling the buildup of large-diameter pines (Taylor et al.,, 2006). Indeed, a comparative analysis of bark beetles found that the best predictor of outbreaks is the availability of susceptible host trees (Koricheva et al.,, 2012). Mountain pine beetle was far more destructive in British Columbia than the western United States (Meddens et al.,, 2012). Although this spatial pattern is not well-researched, we suspect that it involves the high density and spatial continuity of lodgepole pine in the Chilcotin Plateau –– the epicenter of the last two outbreaks in British Columbia.
Our approach differs from previous research on bark beetle dynamics in several key aspects. Rather than focusing on the (proximal) environmental causes of outbreaks, we focus on the (ultimate) beetle life-history causes of outbreaks. Wee present quantitative and empirically-calibrated models that offer more precision than the verbal descriptions found in other reviews of bark beetle dynamics (e.g. Coulson et al.,, 1985; Raffa et al.,, 2008; Weed et al.,, 2015; Biedermann et al.,, 2019; Lantschner and Corley,, 2023; Singh et al.,, 2024) and allow us to identify a minimal set of processes that generate characteristic boom-bust dynamics. Other studies have attempted to link life-history predictors to outbreaking behavior with statistical comparisons (e.g., Nothnagle and Schultz,, 1987; Koricheva et al.,, 2012; Kozlov et al.,, 2010), but this approach faces significant limitations. It’s often unclear how certain predictors are mechanistically connected to outbreaks. For example, larger bark beetles are more prone to outbreaks, contradicting $r-K$ selection theory’s prediction that smaller beetles should have higher growth rates and thus greater outbreak potential (Koricheva et al.,, 2012). More critically, there is very little information available about the biology of non-outbreaking species (Nothnagle and Schultz,, 1987; Johns et al.,, 2016). To circumvent these difficulties, we build a mechanistic model of the bark beetle species for which the most information is available — the mountain pine beetle — and then extrapolate to other species based on shared life history .
The study of excitable dynamics in bark beetles has significant practical implications. The positive feedback mechanisms inherent in excitable systems mean that outbreaks can extend far beyond the spatial extent of an initial disturbance, making them widespread and difficult to manage. This suggests the existence of a “Goldilocks zone” of population densities where management efforts will be most effective –– too low, and the strong Allee effect naturally limits population growth; too high, and the outbreak becomes unmanageable (Hodge et al.,, 2017, Fig. 6). After a major outbreak occurs, there is a known window of time — a ticking clock — before the next outbreak could occur. This period offers forest managers a timeline for implementing preventive measures, such as creating diverse stands in terms of age structure and species composition (Dymond et al.,, 2014; Burton,, 2010).
@@ 1.1 The biology of mountain pine beetle and lodgepole pine
Adult beetles emerge in July or august and After emergence these beetles fly in search of new host trees, typically targeting large pines during the epidemic population phase. Mountain pine beetle has infested Jack pine and Rocky Mountain bristlecone pine, but the general suitability of these species is a matter of ongoing research (Bentz,, 2019; Bentz et al.,, 2022; Bleiker et al.,, 2023). Individual beetles may fly around for several days, but few beetles survive more than 5–10 days outside of the tree (Reid and Reid,, 2008; Fuchs and Borden,, 1985). Emergence from natal trees is synchronous due to temperature-dependent development thresholds (Powell and Bentz,, 2009; Logan and Powell,, 2001), but the synchronicity is not perfect, and so the population-level flight period may last up to 6 weeks (Bleiker and Van Hezewijk,, 2016).
@@
Because individual beetles cannot be tracked, the dispersal phase of the life cycle is mildly mysterious. Available evidence points to three modes of dispersal: short-, medium-, and long-range dispersal. Most beetles disperse less than 30-50 meters under the canopy (Safranyik et al.,, 1992; Robertson et al.,, 2007; for the Douglas-fir beetle Dodds and Ross,, 2002). A minority of beetles disperse up to 5 km (Robertson et al.,, 2009; Powell and Bentz,, 2014; Howe et al.,, 2021; Koch et al.,, 2021), performing pioneer attacks and aggregating to form semi-regular patterns of “spot infestations” (clusters of infested trees) across the landscape (Strohm et al.,, 2013). The smallest fraction of beetles (0–2.5%) fly above the canopy (Robertson et al.,, 2009; Safranyik et al.,, 1992), where they may get carried up to 300 km by upper atmospheric winds (Hiratsuka et al.,, 1982; Cerezke,, 1989; Jackson et al.,, 2008). Some experts have speculated that the fraction may be higher in stands where most suitable host trees have already been killed (Carroll et al.,, 2003; de la Giroday et al.,, 2011; Bleiker,, 2019).
@@
Upon locating a suitable host, MPB females initiate attacks. The trees fight back by exuding resin, which serves to entrap or pitch out (i.e. expel) attacking MPB. Female pioneers release trans and cis-verbenol, aggregation pheromones that draw more beetles to the site (Pureswaran et al.,, 2000; Seybold et al.,, 2018). As males arrive, they emit exo-brevicomin, another aggregation pheromone that specifically attracts additional females (Rudinsky et al.,, 1974). Many bark beetle species use aggregation pheromones and host volatiles to coordinate mass attacks, where swarms of beetles overwhelm a single tree. Depending on the beetle population density, tree suitability, and the tree’s defensive capability, successful mass attacks may last from a couple of hours to a week, continuing until the tree’s resin supply is depleted.
Following a successful mass attack, females bore into the tree and excavate egg galleries. Mating occurs within the tree, and males will either seek out additional females or “settle down” to help with gallery defense. Eggs hatch within one to two weeks, and larvae immediately start feeding on the tree’s phloem. The beetles carry fungal spores in specialized mouthparts, thereby spreading and inoculating the phloem tissue with fungi. Beetle larvae and adults rely on fungal hyphae and spores for nutrition (Reid,, 1958; Bleiker and Six,, 2007). Tree death is hastened by MPB’s fungal symbionts, blue stain fungi spp., which colonize the tree’s transport vessels (Arango-Velez et al.,, 2016). Larvae mine horizontal galleries that terminate in an oval pupal chamber which is lined with spores.
Mature larvae overwinter after accumulating cyroprotectants that allow them to survive temperatures as low as -30C (Bentz and Mullins,, 1999). Other life stages are more susceptible to cold mortality particularly to cold snaps that are early or late in their overwintering period (Wygant,, 1940; Safranyik and Linton,, 1991; Logan et al.,, 1995). Other life stages are more susceptible to cold mortality particularly to cold snaps that are early or late in their overwintering period (Régnière and Bentz,, 2007; Rosenberger et al.,, 2017). Come spring, surviving larvae resume feeding and pupate in the late spring or early summer. By the time of adult emergence, approximately one year after the initial attack, the needles of successfully-attacked trees have turned a rust-red color. This telltale diagnostic makes it possible to gather extensive data through aerial surveys.
@@
Lodgepole pine and ponderosa pine are the predominant hosts of MPB, simply because they are the most abundant pine species in MPB’s geographic range. Lodgepole pine dominates in British Columbia, western Alberta, and the American Rocky Mountain states, whereas ponderosa pine is more abundant in the American southwest and the Sierra mountains (Beaudoin et al.,, 2014; Ellenwood et al.,, 2015). Both species are shade-intolerant and fire-resistant. They tend to be seral (i.e., persisting as a non-climax species in a fire cycle), taking advantage of high-light conditions following a stand-leveling fire (Fryer,, 2018; Anderson,, 2003; Lotan et al.,, 1985). Lodgepole pine has closed cones with seeds that remain viable for decades, effectively creating a seed bank (Teste et al., 2011a, ; Teste et al., 2011b, ). Due to its abundant seeds and its ability to thrive in high-light conditions, lodgepole pine can take advantage of the light gaps created in the years following a mountain pine beetle attack, resulting in the recruitment of seedlings into approximately even-aged cohorts (Axelson et al.,, 2010). However, this regeneration pattern is not universal across MPB’s range; in Alberta, for example, the density of standing dead trees following MPB outbreaks prevents sufficient light penetration, leading to the replacement of lodgepole pine by more shade-tolerant species (Axelson et al.,, 2009).
@@
During epidemics, MPB seeks out large healthy trees with abundant phloem resources. However, in their low-density endemic state, there are not enough beetles to mass attack a single healthy tree, and thus there is a collective behavioral switch where beetles seek out trees that are weakened or stressed, typically due to root disease, wind damage, or lightning (Tkacz and Schmitz,, 1986; Carroll et al., 2006b, ). Endemic populations tend to remain at low density, primarily due to resource limitation, but also because of natural enemies and competition with other bark and woodboring beetle species (Amman,, 1983; Safranyik and Carroll,, 2006). Endemic populations of MPB can transition into an incipient-epidemic state when they successfully breed within trees that have thick phloem. This scenario can occur if a healthy tree experiences acute stress (such as wind damage or lightning) right before the MPB flight period, such that pulse-driven bark beetles have not yet extensively colonized the tree (Bleiker et al.,, 2014). Such stressors can trigger local outbreaks (on the scale of square kilometers). However, widespread outbreaks are usually preceded by widespread drought conditions, which synchronously weaken the resin defenses of many trees (Alfaro et al.,, 2009; Creeden et al.,, 2014; Kolb et al.,, 2016).
1.2 Mountain pine beetle life history